OpenStudy (anonymous):

HELP PLEASE! What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 16 y = x + 4 A. -4 and 0 B. 4 and 0 C. -2 and -3 D. 2 and -3

OpenStudy (anonymous):

@tHe_FiZiCx99 @amistre64 ?

OpenStudy (amistre64):

we have a circle and a line ....

OpenStudy (amistre64):

let y = x+4 in the circle and you create the usual quadratic expression to play with

OpenStudy (anonymous):

a little lost :( I'm sorry

OpenStudy (amistre64):

x2 + y2 = 16 ; this is a circle, at least an equation of one y = x + 4 ; this has to hold for y in any equation in the system, replace it in the circle x2 + (x+4)2 = 16 expand it out and you create a quadratic: ax^2 + bx + c = k, which has plenty of methods to find a solution. OR, just trial and error the options to see what fits

OpenStudy (anonymous):

:l

OpenStudy (anonymous):

I just don't understand any of this

OpenStudy (amistre64):

x2 + (x+4)2 = 16 use this to test the options Does x=0 work? Does x=4 work? Does x=-4 work? etc ....

OpenStudy (anonymous):

x=4 works right? :D

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

would it be B?

OpenStudy (anonymous):

@Mertsj Can you help?

OpenStudy (mertsj):

In the first equation, replace y with x+4

OpenStudy (mertsj):

Now you have only one variable. Solve the equation for x. You should get 2 answers and those are the answers to the question the problem asks.

OpenStudy (anonymous):

okay, can you hang around to make sure I'm right?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

how do I figure out what x^2 is?

OpenStudy (mertsj):

by solving that equation. x^2+(x+4)^2=16

OpenStudy (anonymous):

this is what i have so far after solving for (x+4)^2 x^2 + 8 = 16

OpenStudy (anonymous):

is that good so far?

OpenStudy (mertsj):

(x+4)(x+4)=x^2+8x+16

OpenStudy (anonymous):

combine like terms next?

OpenStudy (mertsj):

yes and get the equation equal to 0 by subtracting 16 from both sides.

OpenStudy (anonymous):

so when i combine like terms does it look like this? 8x^2 = 16

OpenStudy (anonymous):

I think i did something wrong...?

OpenStudy (mertsj):

What is x^2+x^2?

OpenStudy (anonymous):

2x^2?

OpenStudy (mertsj):

Do you know that only like terms can be added?

OpenStudy (anonymous):

yes, I do

OpenStudy (anonymous):

So, what do I do?

OpenStudy (mertsj):

Add the like terms

OpenStudy (mertsj):

How did you get 8x^2 if you know that only like terms can be added?

OpenStudy (anonymous):

I read it wrong. hold on

OpenStudy (anonymous):

I keep confusing myself :l

OpenStudy (mertsj):

Here is what you should have: x^2+(x+4)^2=16 x^2+x^2+8x+16=16

OpenStudy (mertsj):

Now: Add the like terms!!!

OpenStudy (anonymous):

12x + 16 = 16 is what I have so far

OpenStudy (anonymous):

wait!

OpenStudy (mertsj):

Do you know what like terms are?

OpenStudy (anonymous):

x^2, x^2 and 8x are all like terms

OpenStudy (anonymous):

correct?

OpenStudy (mertsj):

Like terms are terms that have the same variables and those variables have the same exponents

OpenStudy (anonymous):

so x^2 and x^2 are like terms

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

when combined i get 2x^4?

OpenStudy (mertsj):

No. 1 cat + 1 cat is 2 cats And 1x^2+1x^2 is 2x^2

OpenStudy (anonymous):

oh, i thought you had to combine the exponents too

OpenStudy (mertsj):

Did you skip Algebra I or something?

OpenStudy (anonymous):

Its been awhile....i did that freshmen year and now im a senior

OpenStudy (anonymous):

I don't retain math very well as you can see...

OpenStudy (anonymous):

I think I'll just figure it out on my own...

OpenStudy (mertsj):

So now you have 2x^2+8x+16=16 right?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Subtract 16 from both sides.

OpenStudy (anonymous):

so now I combine 16 with 16 correct?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now i have 2x^2 + 8x = 0

OpenStudy (mertsj):

Now divide both sides by 2

OpenStudy (anonymous):

x^2 + 4x = 0

OpenStudy (mertsj):

Now factor out x

OpenStudy (anonymous):

4x^2?

OpenStudy (mertsj):

x^2+4x=x(x+4)

OpenStudy (mertsj):

Now set each factor equal to 0 and solve each equation.

OpenStudy (anonymous):

so it would be B?

OpenStudy (mertsj):

no

OpenStudy (anonymous):

darn

OpenStudy (mertsj):

Did you set each factor equal to 0?